# EquiReg: Equivariance Regularized Diffusion for Inverse Problems

### v1

@@ Appendix A EquiReg for Text-to-Image Guidance
Given the “source” image, DreamSampler [82] is asked to transform the source image using the prompt. Applying EquiReg to DreamSampler, we observe perceptual improvement of generated images as well as artifact reduction. Figure 7 shows a cat, as the source image, being transformed into “corgi”, “croissant”, “cat statue”, and “fluffy dog”. Equi-DreamSampler generates more realistic images than DreamSampler (e.g., EquiReg modifies the three front legged corgi to two front legs).
Figure 7: EquiReg helps the text-guided DreamSampler generate more realistic images.
Figure 8: Impact of EquiReg parameter $\lambda$, implicit acceleration, and introduction of more image details. Women, red hair, smiling.
We have also observed an implicit acceleration of image generation when EquiReg is imposed (Figure 8). Equi-DreamSampler with $50$ DDIM steps can generate images that are only possible with DreamSampler when the DDIM steps are increased. We attribute this to EquiReg’s ability to generate images that are closer to the data manifold. For example, the increase of DDIM steps in DreamSampler (from $50$ to $75$ to $100$) has a relatively similar effect to increasing the EquiReg regularizer $\lambda$ at a fixed $50$ DDIM steps. Figure 8 shows that increasing the regularization $\lambda$ results in addition of a car in the background. For DreamSampler, an early notion of the car seat in the background start to arise only when DDIM is increased to $100$ (see also Figures 9 and 10).
Figure 9: Adding EquiReg into the text-to-image guidance method DreamSampler for improved performance. Boy, curly hair.
@@
Appendix B Additional Experiments on Image Restoration Tasks
We show two additional results. EquiReg improves LatentDAPS performance on ImageNet for box inpainting and motion deblur (Table 5). Moreover, Equi-DPS performs much better than DPS, increasing the PSNR from $20.70$ to $23.45$ dB for super-resolution on ImageNet (Table 6).
Table 5: Equi-LatentDAPS vs. LatentDAPS on ImageNet 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.05$.
Method Box inpainting Motion deblur
@@ Appendix B Additional Experiments on Image Restoration Tasks
LatentDAPS 0.310 20.19 0.305 25.08
Equi-LatentDAPS (ours) 0.299 20.67 0.286 25.86
Table 6: Equi-DPS vs. DPS on ImageNet 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.05$.
Method Super-resolution (x4) Random inpainting
[…]
DPS 169.6 20.70 94.83 25.96
Equi-DPS (ours) 102.7 23.45 55.87 28.32
Appendix C Additional Information on Image Restoration Tasks
We report standard deviation on all linear tasks for PSLD, Equi-PSLD, and EquiCon-PSLD (Table 7).
Table 7: Reporting standard deviations. Inverse problems on FFHQ 256 $\times$ 256 with $\sigma_{y}=0.05$.
[…]
Gaussian deblur PSLD 0.357 (0.060) 22.87 (1.67) 0.537 (0.094)
Equi-PSLD (ours) 0.344 (0.055) 24.42 (1.66) 0.657 (0.076)
EquiCon-PSLD (ours) 0.320 (0.048) 24.38 (1.55) 0.664 (0.068)
Motion deblur PSLD 0.322 (0.048) 24.25 (1.47) 0.615 (0.075)
Equi-PSLD (ours) 0.338 (0.052) 24.83 (1.75) 0.669 (0.068)
EquiCon-PSLD (ours) 0.322 (0.054) 25.14 (1.62) 0.699 (0.064)
Super-res. (×$\times 4$) PSLD 0.313 (0.098) 24.51 (2.98) 0.650 (0.140)
Equi-PSLD (ours) 0.289 (0.088) 26.32 (2.24) 0.718 (0.104)
EquiCon-PSLD (ours) 0.277 (0.093) 26.14 (2.74) 0.722 (0.119)
Box inpainting PSLD 0.158 (0.023) 24.22 (2.80) 0.819 (0.031)
Equi-PSLD (ours) 0.098 (0.019) 24.19 (2.67) 0.875 (0.027)
EquiCon-PSLD (ours) 0.092 (0.021) 24.26 (2.80) 0.884 (0.029)
Random inpainting PSLD 0.246 (0.053) 29.05 (1.98) 0.809 (0.049)
Equi-PSLD (ours) 0.188 (0.034) 30.43 (1.76) 0.878 (0.027)
EquiCon-PSLD (ours) 0.204 (0.051) 29.99 (2.06) 0.847 (0.047)
Appendix D Visualizations for Image Restoration Experiments
@@
(b) Example visualizations of used images.
Figure 11: Training induced equivariance for a pre-trained encoder. Experimental setup corresponds to Figure 4(a) in the main text, which uses a pre-trained decoder.
Figure 12: Equivariance error of a pre-trained decoder over different Gaussian noise levels $\sigma$. We apply different Gaussian noise levels to clean images to see the effect of noise on equivariance error. We observe that as Gaussian noise increases, equivariance error increases steadily.
@@
(d) Super-resolution (×$\times 4$).
@@
(b) Super-resolution (×$\times 4$).
@@ Appendix E Implementation Details for Image Restoration Tasks
Hyperparameters. Our method introduces a single hyperparameter $\lambda$ that controls the amount of regularization applied. Below we include a table detailing the use of this hyperparameter in the main experiments (Table 8). For majority of experiments, we keep $\lambda$ constant throughout iterations. However, we note that in our Equi-DPS and Equi-LatentDAPS experiments, we scale $\lambda$ by the noise scheduler at every iteration. For all unscaled experiments, we employ early stopping, setting $\lambda=0$ for the last 10% of sampling.
Table 8: Equivariance regularization weight $\lambda$ used in main experiments.
Method Box Inpainting Random Inpainting Gaussian Deblur Motion Deblur Super-resolution (×$\times 4$)
@@
PSLD. We integrate EquiReg into PSLD by simply adding an additional gradient update step using our regularization term (Algorithm 2, Algorithm 3).
In our experiments, we use the official PSLD implementation from Rout et al. [38], running with its default settings to reproduce the baseline results. We note that in our code, we do not square the norm when computing the gradient, aligning with PSLD’s implementation.
@@
1:$T,\bm{y},\{\eta_{t}\}_{t=1}^{T},\{\gamma_{t}\}_{t=1}^{T},\{\tilde{\sigma}_{t}\}_{t=1}^{T}$
2:$\mathcal{E},\mathcal{D},\mathcal{A}{\bm{x}}_{0}^{\ast},\mathcal{A},\bm{s}_{\theta},$ $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
3:${\bm{z}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
4:for $t=T-1$ to 00 do
@@
1:$T,\bm{y},\{\eta_{t}\}_{t=1}^{T},\{\gamma_{t}\}_{t=1}^{T},\{\tilde{\sigma}_{t}\}_{t=1}^{T}$
2:$\mathcal{E},\mathcal{D},\mathcal{A}{\bm{x}}_{0}^{\ast},\mathcal{A},\bm{s}_{\theta},$ $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
3:${\bm{z}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
4:for $t=T-1$ to 00 do
@@
ReSample. We integrate EquiReg into ReSample by adding our regularization term into the hard data consistency step (Algorithm 4, Algorithm 5). We note that the ReSample algorithm employs a two-stage approach; initially, it performs pixel-space optimization, and later it performs latent-space optimization. We apply EquiReg in the latent-space optimization stage.
In our experiments, we use the official ReSample implementation from Song et al. [37], running with its default settings to reproduce the baseline results.
@@
2:${\bm{z}}_{T}\sim\mathcal{N}(\mathbf{0},\bm{I})$triangleright Initial noise vector
@@
5:    $\hat{\bm{\epsilon}}_{t+1}={\bm{s}}_{\theta}({\bm{z}}_{t+1},t+1)$triangleright Compute the score
6:    $\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})=\frac{1}{\sqrt{\bar{\alpha}_{t+1}}}({\bm{z}}_{t+1}-\sqrt{1-\bar{\alpha}_{t+1}}\hat{\bm{\epsilon}}_{t+1})$triangleright Predict $\hat{{\bm{z}}}_{0}$ using Tweedie’s formula
7:    ${\bm{z}}^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})+\sqrt{1-\bar{\alpha}_{t}-\eta\delta^{2}}\hat{\bm{\epsilon}}_{t+1}+\eta\delta\bm{\epsilon}_{1}$triangleright Unconditional DDIM step
@@
14:         ${\bm{z}}_{t}=\text{StochasticResample}(\hat{{\bm{z}}}_{0}({\bm{y}}),{\bm{z}}^{\prime}_{t},\gamma)$triangleright Map back to $t$
@@
16:         ${\bm{z}}_{t}={\bm{z}}^{\prime}_{t}$triangleright Unconditional sampling if not resampling
@@
2:${\bm{z}}_{T}\sim\mathcal{N}(\mathbf{0},\bm{I})$triangleright Initial noise vector
@@
5:    $\hat{\bm{\epsilon}}_{t+1}={\bm{s}}_{\theta}({\bm{z}}_{t+1},t+1)$triangleright Compute the score
6:    $\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})=\frac{1}{\sqrt{\bar{\alpha}_{t+1}}}({\bm{z}}_{t+1}-\sqrt{1-\bar{\alpha}_{t+1}}\hat{\bm{\epsilon}}_{t+1})$triangleright Predict $\hat{{\bm{z}}}_{0}$ using Tweedie’s formula
7:    ${\bm{z}}^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})+\sqrt{1-\bar{\alpha}_{t}-\eta\delta^{2}}\hat{\bm{\epsilon}}_{t+1}+\eta\delta\bm{\epsilon}_{1}$triangleright Unconditional DDIM step
@@
14:         ${\bm{z}}_{t}=\text{StochasticResample}(\hat{{\bm{z}}}_{0}({\bm{y}}),{\bm{z}}^{\prime}_{t},\gamma)$triangleright Map back to $t$
@@
16:         ${\bm{z}}_{t}={\bm{z}}^{\prime}_{t}$triangleright Unconditional sampling if not resampling
@@
DPS. Similar to PSLD, we integrate EquiReg into DPS by simply adding an additional gradient update step using our regularization term (Algorithm 6).
In our experiments, we use the official DPS implementation from Chung et al. [32], running with its default settings to reproduce the baseline results.
@@
LatentDAPS. We integrate EquiReg into LatentDAPS by simply adding an additional gradient update step using our regularization term (Algorithm 7).
In our experiments, we use the official LatentDAPS implementation from Zhang et al. [40], running with its default settings to reproduce the baseline results.
Algorithm 7 Equi-LatentDAPS for Image Restoration Tasks
1:Latent space score model ${\bm{s}}_{\theta}$, measurement ${\bm{y}}$, noise schedule $\sigma_{t}$, $\{t_{i}\}_{i\in\{0,\ldots,N_{A}\}}$, encoder $\mathcal{E}$ and decoder $\mathcal{D}$,  $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
2:Sample ${\bm{z}}_{T}\sim\mathcal{N}(\mathbf{0},\sigma_{T}^{2}\bm{I})$
3:for $i=N_{A},N_{A}-1,\ldots,1$ do
4:    Compute $\hat{{\bm{z}}}_{0}^{(0)}=\hat{{\bm{z}}}_{0}({\bm{z}}_{t_{i}})$ by solving the probability flow ODE with ${\bm{s}}_{\theta}$
5:    for $j=0,\ldots,N-1$ do
6:          $\bm{g}\leftarrow\nabla_{\hat{{\bm{z}}}_{0}}\log p(\hat{{\bm{z}}}_{0}^{(j)}|{\bm{z}}_{t_{i}})+\nabla_{\hat{{\bm{z}}}_{0}}\log p({\bm{y}}|\hat{{\bm{z}}}_{0}^{(j)})+\lambda_{t}\nabla_{\hat{{\bm{z}}}_{0}}\|S_{g}(\mathcal{D}(\hat{{\bm{z}}}_{0}^{(j)}))-\mathcal{D}(T_{g}(\hat{{\bm{z}}}_{0}^{(j)}))\|_{2}^{2}$
7:         $\hat{{\bm{z}}}_{0}^{(j+1)}\leftarrow\hat{{\bm{z}}}_{0}^{(j)}+\eta_{t}\bm{g}+\sqrt{2\eta_{t}}\bm{\epsilon}_{j}$, $\bm{\epsilon}_{j}\sim\mathcal{N}(\mathbf{0},\bm{I})$
8:    end for
9:    Sample ${\bm{z}}_{t_{i-1}}\sim\mathcal{N}(\hat{{\bm{z}}}_{0}^{(N)},\sigma_{t_{i-1}}^{2}\bm{I})$
10:end for
11:return $\mathcal{D}({\bm{z}}_{0})$

### v2

@@ Appendix A EquiReg for Text-to-Image Guidance
Figure 7: Impact of EquiReg parameter $\lambda_{t}$, implicit acceleration, and introduction of more image details on FFHQ $\times512$. Women, red hair, smiling.
Figure 8: Adding EquiReg into the text-to-image guidance method DreamSampler for improved performance on FFHQ $\times512$. Boy, curly hair.
@@
Figure 10: Adding EquiReg into the text-to-image guidance method DreamSampler for improved performance. Mustache, beard.
Figure 11: Adding EquiReg into the text-to-image guidance method DreamSampler for improved performance. Sculpture.
Appendix B Additional Experiments on Robustness
Figure 12: Advantages of EquiReg under reduced DDIM steps. Super-resolution on FFHQ.
Table 9: EquiReg improves SITCOM under reduced measurement consistency steps ($K_{\text{meas}}$). Motion deblur on FFHQ sampled with $50$ DDIM steps.
@@ Appendix B Additional Experiments on Robustness
10 N/A 28.06 0.81 21.57
10 1 28.71 0.82 21.07
5 5 29.26 0.83 11.09
20 N/A 27.04 0.79 38.85
20 1 28.54 0.82 37.74
10 10 28.93 0.82 20.92
30 N/A 27.79 0.80 58.84
30 1 28.35 0.81 55.51
15 15 29.63 0.84 30.19
40 N/A 30.40 0.85 78.08
40 1 30.58 0.85 69.83
20 20 29.50 0.83 41.02
60 N/A 28.35 0.81 108.57
60 1 27.02 0.78 95.62
30 30 31.36 0.87 59.38
Table 10: EquiReg Effectiveness with Subset of Group Actions.
PSLD Equi-PSLD (90, 270 deg)
$uparrow PSNR\uparrow SSIM\uparrow$
15.86 (1.19) 0.77 (0.03) 17.60 (1.60) 0.79 (0.03)
Appendix C Visualizations for Image Restoration Experiments
@@
(b) Example visualizations of used images and corresponding equivariance error computed using the decoder (see Figure˜4(a)).
Figure 14: Training induced equivariance for a pre-trained function.
@@
(d) Super-resolution ($\times 4$).
@@
(b) Super-resolution ($\times 4$).
@@
Appendix D Diversity Analysis
In the Bayesian setting, the objective of solving inverse problems with diffusion models is to sample from high-probability regions of the posterior distribution. While the goal is not to maximize “diversity”, the true diversity emerges when the posterior admits meaningful variability. In practice, diversity-related concerns in inverse problems arise when a method suffers from mode collapse, i.e., the sampler becomes biased and fails to explore multiple plausible modes of the posterior. Thus, the relevant question is whether a method properly explores the posterior rather than whether it maximizes diversity in an unconstrained sense.
Because closed-form posteriors are unavailable for real image restoration tasks, the standard practice in the diffusion inverse-problem literature is to evaluate diversity through variation among plausible reconstructions consistent with the measurement, without collapsing to a single solution. This is the notion of “diversity” our work adopts.
Given the goal of posterior sampling, EquiReg is not designed to maximize diversity for its own sake. Its objective is to incorporate data-inherent geometric structure (equivariance) to guide sampling toward high-probability regions of the posterior. Hence, diversity arises naturally from the ill-posedness of the inverse problem; it is a consequence of posterior uncertainty, not the goal of the regularizer.
To quantify this effect, in addition to reconstruction quality, we analyzed the diversity of posterior samples produced by EquiReg. We evaluate diversity metrics across multiple tasks and difficulty levels to characterize the sampling behavior of our method.
D.1 Experimental Setup
[…]
D.2 Results and Discussion
Table 7 shows that Equi-DPS achieves favorable fidelity-diversity trade-offs across three inverse problems. For box inpainting and super-resolution, equivariance regularization improves both fidelity and diversity simultaneously. For Gaussian deblurring, Equi-DPS achieves 15-20% better fidelity while retaining 80-85% of baseline diversity, representing a modest but justified trade-off. These results demonstrate that equivariance constraints do not inherently suppress diversity; rather, they can guide sampling toward regions of higher data fidelity while maintaining posterior exploration.
Figure 17 reveals linear diversity scaling with task difficulty. Diversity metrics grow proportionally with task difficulty, indicating Equi-DPS naturally expands sampling as problems become more ill-posed. This linear relationship demonstrates stable, predictable behavior across difficulty levels without artificial diversity suppression. Figures 6 and 18 provide qualitative results.
(a) Box inpainting
(b) Gaussian deblur
(c) Super-resolution
Figure 17: Diversity vs task difficulty across three inverse problems. As task difficulty increases (larger inpainting mask, stronger blur, higher SR scale), both diversity metrics increase proportionally, demonstrating that Equi-DPS maintains healthy posterior sampling behavior across a wide difficulty spectrum.
(a) Gaussian deblur
(b) Box inpainting
Figure 18: Qualitative diversity examples across Gaussian deblur and box inpainting. Each subfigure shows $K=4$ posterior samples for two different test images. (a) Gaussian deblur: samples differ in facial expressions and accessories (i.e., earrings in first test image). (b) Box inpainting: Inputs were obstructed with $160\times 160$ masks. Samples exhibit perceptually distinct facial features (i.e., expressions, eye gaze, facial structure). Across both tasks, EquiReg produces diverse plausible reconstructions rather than collapsing to a single mode.
D.3 Conclusion
Finally, we highlight that EquiReg improves both fidelity and diversity on 2 of the 3 considered tasks, an encouraging outcome that is uncommon given the general behavior of classical regularizers. Hand-crafted regularizers such as TV and $\ell_{1}$ may suppress diversity by shrinking solutions toward simple structures. By contrast, EquiReg leverages data-dependent regularization that captures the richness and structural complexity of the underlying data manifold, enabling it to preserve manifold-consistent variability while suppressing implausible samples.
High diversity without fidelity is not meaningful for posterior sampling. A method that samples the entire solution space, including low-probability and artifacted regions, may score well on diversity but fail to provide useful reconstructions. Equi-DPS avoids this failure mode: it maintains meaningful diversity while reducing artifacts and improving perceptual quality. In the experiments conducted during the rebuttal, our goal was to demonstrate clearly that EquiReg preserves meaningful diversity, reflecting the posterior uncertainty, rather than unstructured or unconstrained variability.
@@ Appendix E Implementation Details for Image Restoration Tasks
Experimental Setup. We evaluate EquiReg on a variety of linear and nonlinear restoration tasks for natural images. We fix sets of $100$ images from FFHQ and ImageNet as our validation sets. All images are normalized from $[0,1]$. For the majority of experiments, we use noise level $\sigma_{{\bm{y}}}=0.05$ (we indicate $\sigma_{{\bm{y}}}$ in our tables). For linear inverse problems, we consider (1) box inpainting, (2) random inpainting, (3) Gaussian deblur, (4) motion deblur, and (5) super-resolution. We apply a random $128\times 128$ pixel box for box inpainting, and a $70\%$ random mask for random inpainting. For Gaussian and motion deblur, we use kernels of size $61\times 61$, with standard deviations of $3.0$ and $0.5$, respectively. For super-resolution, we downscale images by a factor of $4$ using a bicubic resizer. For nonlinear inverse problems, we consider (1) phase retrieval, (2) nonlinear deblur, and (3) high dynamic range (HDR). We use an oversampling rate of $2.0$ for phase retrieval, and due to instability of the task, we generate four independent reconstructions and take the best result (as also done in DPS (Chung et al., 2023), DAPS (Zhang et al., 2025a), and DiffStateGrad (Zirvi et al., 2025)). We use the default setting from (Tran et al., 2021) for nonlinear deblur, and a scale factor of 2 for HDR.
Hyperparameters. Our method introduces a single hyperparameter $\lambda_{t}$ that controls the amount of regularization applied. Below we include a table detailing the use of this hyperparameter in the main experiments (Table˜11). For majority of experiments, we keep $\lambda_{t}$ constant throughout iterations. For all unscaled experiments, we employ early stopping, setting $\lambda_{t}=0$ for the last 10% of sampling.
Table 11: Equivariance regularization weight $\lambda_{t}$ used in main experiments.
Method Box Inpainting Random Inpainting Gaussian Deblur Motion Deblur Super-resolution ($\times 4$)
@@
[…]
4:for t=T1t=T-1 to 0 do
@@
PSLD. We integrate EquiReg into PSLD by simply adding an additional gradient update step using our regularization term (Algorithms˜2 and 3).
In our experiments, we use the official PSLD implementation from Rout et al. (2023), running with its default settings to reproduce the baseline results. We note that in our code, we do not square the norm when computing the gradient, aligning with PSLD’s implementation.
@@
[…]
4:for t=T1t=T-1 to 0 do
@@
ReSample. We integrate EquiReg into ReSample by adding our regularization term into the hard data consistency step (Algorithms˜4 and 5). We note that the ReSample algorithm employs a two-stage approach; initially, it performs pixel-space optimization, and later it performs latent-space optimization. We apply EquiReg in the latent-space optimization stage.
In our experiments, we use the official ReSample implementation from Song et al. (2023a), running with its default settings to reproduce the baseline results.
@@
[…]
@@
5:  $\hat{\bm{\epsilon}}_{t+1}={\bm{s}}_{\theta}({\bm{z}}_{t+1},t+1)$ $\triangleright$ Compute the score
6:  $\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})=\frac{1}{\sqrt{\bar{\alpha}_{t+1}}}({\bm{z}}_{t+1}-\sqrt{1-\bar{\alpha}_{t+1}}\hat{\bm{\epsilon}}_{t+1})$ $\triangleright$ Predict $\hat{{\bm{z}}}_{0}$ using Tweedie’s formula
7:  ${\bm{z}}^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})+\sqrt{1-\bar{\alpha}_{t}-\eta\delta^{2}}\hat{\bm{\epsilon}}_{t+1}+\eta\delta\bm{\epsilon}_{1}$ $\triangleright$ Unconditional DDIM step
@@
14:    ${\bm{z}}_{t}=\text{StochasticResample}(\hat{{\bm{z}}}_{0}({\bm{y}}),{\bm{z}}^{\prime}_{t},\gamma)$ $\triangleright$ Map back to $t$
@@
16:    ${\bm{z}}_{t}={\bm{z}}^{\prime}_{t}$ $\triangleright$ Unconditional sampling if not resampling
@@
[…]
@@
5:  $\hat{\bm{\epsilon}}_{t+1}={\bm{s}}_{\theta}({\bm{z}}_{t+1},t+1)$ $\triangleright$ Compute the score
6:  $\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})=\frac{1}{\sqrt{\bar{\alpha}_{t+1}}}({\bm{z}}_{t+1}-\sqrt{1-\bar{\alpha}_{t+1}}\hat{\bm{\epsilon}}_{t+1})$ $\triangleright$ Predict $\hat{{\bm{z}}}_{0}$ using Tweedie’s formula
7:  ${\bm{z}}^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})+\sqrt{1-\bar{\alpha}_{t}-\eta\delta^{2}}\hat{\bm{\epsilon}}_{t+1}+\eta\delta\bm{\epsilon}_{1}$ $\triangleright$ Unconditional DDIM step
@@
14:    ${\bm{z}}_{t}=\text{StochasticResample}(\hat{{\bm{z}}}_{0}({\bm{y}}),{\bm{z}}^{\prime}_{t},\gamma)$ $\triangleright$ Map back to $t$
@@
16:    ${\bm{z}}_{t}={\bm{z}}^{\prime}_{t}$ $\triangleright$ Unconditional sampling if not resampling
@@
DPS. Similar to PSLD, we integrate EquiReg into DPS by simply adding an additional gradient update step using our regularization term (Algorithm˜6).
In our experiments, we use the official DPS implementation from Chung et al. (2023), running with its default settings to reproduce the baseline results.
@@
SITCOM. We augment the original SITCOM algorithm by introducing an additional equivariant refinement stage at each reverse diffusion step. After completing the standard measurement and backward-consistency gradient updates, we perform a second optimization over the equivariance loss, enforcing consistency between $\mathcal{E}(T_{g}(v))$ and $T_{g}(\mathcal{E}(v))$ (Algorithm˜7).
In our experiments, we use the official SITCOM implementation from Alkhouri et al. (2025), running with its default settings to reproduce the baseline results.
Algorithm 7 Equi-SITCOM for Image Restoration Tasks
1:Measurements $\mathbf{y}$, forward operator $\mathcal{A}(\cdot)$, pre-trained DM $\epsilon_{\theta}(\cdot,\cdot)$, diffusion steps NN, schedule $\bar{\alpha}_{i}$, measurement gradient steps $K$, equivariant gradient steps $K_{\text{equi}}$, stop $\delta$, lr $\gamma$, reg. $\lambda$.
2:Restored image $\hat{\mathbf{x}}$.
3:Initialize $\mathbf{x}_{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $\Delta t=\left\lfloor\tfrac{T}{N}\right\rfloor$.
4:for $i=N,N-1,\ldots,1$ do $\triangleright$ Reducing diffusion sampling steps
5:  $\mathbf{v}_{i}^{(0)}\leftarrow\mathbf{x}_{i}$ $\triangleright$ Init for closeness (C3)
6:  for $k=1,\ldots,K$ do $\triangleright$ Adam on measurement/backward consistency (C1, C2)
7:    $\mathbf{v}_{i}^{(k)}\leftarrow\mathbf{v}_{i}^{(k-1)}-\gamma\nabla_{\mathbf{v}_{i}}\!\left[\left\|\mathcal{A}\!\left(\tfrac{1}{\sqrt{\bar{\alpha}_{i}}}\big(\mathbf{v}_{i}-\sqrt{1-\bar{\alpha}_{i}}\,\epsilon_{\theta}(\mathbf{v}_{i},i\Delta t)\big)\right)-\mathbf{y}\right\|_{2}^{2}+\lambda\|\mathbf{x}_{i}-\mathbf{v}_{i}\|_{2}^{2}\right]\Big|_{\mathbf{v}_{i}=\mathbf{v}_{i}^{(k-1)}}$
8:    if $\left\|\mathcal{A}\!\left(\tfrac{1}{\sqrt{\bar{\alpha}_{i}}}\big(\mathbf{v}_{i}^{(k)}-\sqrt{1-\bar{\alpha}_{i}}\,\epsilon_{\theta}(\mathbf{v}_{i}^{(k)},i\Delta t)\big)\right)-\mathbf{y}\right\|_{2}^{2}<\delta^{2}$ then
9:     break $\triangleright$ Prevent noise overfitting
10:    end if
11:  end for
12:  $\mathbf{v}_{i}^{(0)}\leftarrow\mathbf{v}_{i}^{(k)}$ $\triangleright$ Initialize to optimized $\mathbf{v}_{i}$
13:  for $k=1,\ldots,K_{\text{equi}}$ do
14:    $\mathbf{v}_{i}^{(k)}\leftarrow\mathbf{v}_{i}^{(k-1)}-\gamma\nabla_{\mathbf{v}_{i}}\!\left[\left\|\mathcal{E}\!\big(T_{g}(\mathbf{v}_{i}^{(k)})\big)-T_{g}\!\big(\mathcal{E}(\mathbf{v}_{i}^{(k)})\big)\right\|_{2}^{2}\right]\Big|_{\mathbf{v}_{i}=\mathbf{v}_{i}^{(k-1)}}$
15:    if $\left\|\mathcal{E}\!\big(T_{g}(\mathbf{v}_{i}^{(k)})\big)-T_{g}\!\big(\mathcal{E}(\mathbf{v}_{i}^{(k)})\big)\right\|_{2}^{2}<\delta^{2}$ then
16:     break
17:    end if
18:  end for
19:  $\hat{\mathbf{v}}_{i}\leftarrow\mathbf{v}_{i}^{(k)}$ $\triangleright$ Backward diffusion consistency (C2)
20:  $\hat{\mathbf{x}}^{\prime}_{0}\leftarrow\tfrac{1}{\sqrt{\bar{\alpha}_{i}}}\!\left[\hat{\mathbf{v}}_{i}-\sqrt{1-\bar{\alpha}_{i}}\,\epsilon_{\theta}(\hat{\mathbf{v}}_{i},i\Delta t)\right]$ $\triangleright$ Backward consistency (C2)
21:  $\mathbf{x}_{i-1}\leftarrow\sqrt{\bar{\alpha}_{i-1}}\,\hat{\mathbf{x}}^{\prime}_{0}+\sqrt{1-\bar{\alpha}_{i-1}}\,\bm{\eta}_{i}$, $\bm{\eta}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ $\triangleright$ Forward consistency (C3)
22:end for
23:return $\hat{\mathbf{x}}=\mathbf{x}_{0}$